how u solve nd check 6x-2=2x-1
Answer by lefty4ever26(59)

6x-2=2x-1 Subtract 2x from both sides
-2x -2x
4x-2=-1 Add 2 to both sides
+2 +2
4x=1 Divide by 4 on both sides
/4 /4
x=1/4 To check plug (1/4) back into the original equation wherever there is an x. If the two sides equal it is the right answer.
6(1/4)-2=2(1/4)-1
(3/2)-2=(1/2)-1
(-1/2)=(-1/2)
So x=(1/4) is the right answer.